OpenStudy (anonymous):

integrate x ^(-7) dx

hartnn (hartnn):

use the standard result \(\huge \int x^ndx = \dfrac{x^{n+1}}{n+1}+c\)

hartnn (hartnn):

here, n = -7

OpenStudy (anonymous):

-(1/6) x^-6

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

but isn't it if you integrate 1/x = ln|x|?

OpenStudy (anonymous):

a little confused because the original problem can be written as 1/x^7

hartnn (hartnn):

it can be written, but to be able to use standard formula, you'll have to write 1/x^7 as x^-7 again :)

OpenStudy (anonymous):

ok